Pl. explain Find the missing character from among the alternatives. 13 19 4, 71 9 10, 128 32 ? Ans 1) 10 2) 15 3) 20 4) 25

Pl. explain
Find the missing character from among the alternatives.
 
13   19   4,
71    9  10,
128  32   ?
 
 
Ans
 
1) 10
2) 15
3) 20
4) 25

Grade:11

1 Answers

Shobhit Varshney IIT Roorkee
10 years ago
Dear student,

correct answer is 20.

(13 + 19)/4 = 8
(71+9)/9 = 8
….….….….….….….........................................................................................
Thanks